The force F (in pounds) acting at an angle à Ž ¸ with the horizontal that is needed to drag a crate weighing W pounds
along a horizontal surface at a constant velocity is given by
1
Expert's answer
2021-10-18T15:52:47-0400

The force F (in pounds) acting at an angle "\theta" with the horizontal that is needed to drag a crate weighing W pounds along a horizontal surface at a constant velocity is given by "F=\dfrac{\mu W}{\cos \theta+\mu \sin \theta}" where "\mu" is a constant called the coefficient of sliding friction between the crate and the surface. Suppose that the crate weighs "150 \mathrm{lb}" and that "\mu=0.3"

- Find "d F / d \theta" when "\theta=30^{\circ}" . Express the answer in units of pounds/degree.


"\theta = 30° × \dfracπ{180°}\ rad= \dfracπ{6}\text{ rad}"


"F= \dfrac{0.3× 150}{(\cosθ+0.3\sinθ)}"


"F= \dfrac{45}{(\cosθ+0.3\sinθ)}"


"\dfrac{dF}{dθ}=\dfrac{[(\cosθ+0.3\sinθ)(\dfrac{d}{dθ}45)]-[45(\dfrac d{dθ}\cosθ + 0.3\dfrac d{dθ }\sinθ)] }{(\cosθ+0.3\sinθ)^2 }"


"\dfrac{dF}{dθ}=\dfrac{[-45(-\sinθ+0.3\cosθ)]}{(\cosθ+0.3\sinθ)^2}"


"\dfrac{dF}{dθ} =\dfrac{(45\sinθ-13.5\cosθ)}{(\cosθ+0.3\sinθ)^2}"